The area of a rectangle is 42 ft. squared, , and the length of the rectangle is 11 ft. less than three times the width. Find the dimensions of the rectangle:
Length: ft.
Width: ft.
PLEASE HELP!
THANK YOU :)

Answer by stanbon(75887)

The area of a rectangle is 42 ft. squared, , and the length of the rectangle is 11 ft. less than three times the width. Find the dimensions of the rectangle:
------
Let width be "x"
Then length is "3x-11"
---
Equation:
width*length = 42
x(3x-11) = 42
3x^2-11x-42 = 0
-----
x = [11+-sqrt(121-4*3*-42))]/6
x = [11+-sqrt(625)]/6
Positive solution:
x = (11+25)/6
x = 36/6 = 6 ft (width)
3x-11 = 18-11 = 7 ft (length)
---------------
Cheers,
Stan H.
================